OpenStudy (anonymous):

v(t)=radical(t^7) - radical(t) find the second derivative v"(t).

OpenStudy (anonymous):

\[v(t)=\sqrt{t^7}-\sqrt{t}\] maybe exponential notation will help with second derivative \[v(t)=t^{\frac{7}{2}}-t^{\frac{1}{2}}\] then power rule twice

OpenStudy (anonymous):

first time get \[\frac{7}{2}t^{\frac{5}{2}}-\frac{1}{2}t^{-\frac{1}{2}}\] one more time and you are done

OpenStudy (anonymous):

Thanks Satellite!